OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y − z, if x = 2, y = 3, and z = −6

OpenStudy (anonymous):

same thing. if x = 2 then 4x = 8. so 8-6--6

OpenStudy (lalaly):

4*2 - 2*3 -(-6) = 8-6+6=8

OpenStudy (anonymous):

woah,, wich ones rite?

OpenStudy (anonymous):

they both mean the same thing

OpenStudy (anonymous):

lalaly has shown you what it looks like when you plug the numbers in for x y and z, then she showed you the additon and subtraction

OpenStudy (lalaly):

@mybrainvsme i like ur username xD

OpenStudy (anonymous):

:D cool. It's definitely true.

OpenStudy (anonymous):

me tooo